OpenStudy (anonymous):

find the domain of the function algebraically.....(square root 4-x)/((x+1)(x^2+1)

hero (hero):

\[\frac{(\sqrt{4-x})}{(x+1)(x^2+1)}\]

hero (hero):

Is that it?

OpenStudy (anonymous):

The value under the square root must not be negative.

OpenStudy (anonymous):

Thus x < 4 is required.

OpenStudy (anonymous):

Also, x+1 need not be 0. So x can not be -1.

OpenStudy (anonymous):

from the radical we have\[4-x \ge0\]or\[x \le4\]but from the denominator we have\[x \neq1\]if you put all that together you get\[(-\infty,1) \cup(1,4]\]

OpenStudy (anonymous):

So the domain is \[ (-\infty, 4] \backslash \{ -1 \} \]

OpenStudy (anonymous):

= (-\infty, -1) \cup (-1, 4]

OpenStudy (anonymous):

omg thank you guys soooooo much i just wasted an hour watching youtube videos for no reason =] thanks again!!